Integrals Giving Inverse Trigonometric Functions


∫ dx/√(9-x²)


1
Expert's answer
2021-04-15T07:20:11-0400

We know that ,"\int \dfrac{dx}{\sqrt{a^2-x^2}}=sin^{-1}\dfrac{x}{a}+ C"


"\Rightarrow \int \dfrac{dx}{\sqrt{9-x^2}}=\int \dfrac{dx}{\sqrt{3^2-x^2}}= sin^{-1}\dfrac{x}{3} +C"